1. To derive the 5-year forward rate three years from now, the information required are:

A. 2-year and 3-year spot rates.

B. 3-year and 5-year spot rates.

C. 3-year and 8-year spot rates.

2. For a mortgage pass-through security, which of the following risks most likely increases as interest rates decline?

A. Credit.

B. Extension.

C. Contraction.

3. Suppose an annual coupon payment bond has a maturity of 10 years. If an investor's total return is $210.59 on a bond that is priced at $98.42, what is the realized rate of return?

A. 1.3%

B. 7.9%

C. 114%
Answer #1

Q1) C) 3-year and 8-year spot rates.

explanation: formula for forward rate is as follows:-

(1+forward rate(3,5))^5 = (1+spot rate(8))^8 / (1+spot rate(3))^3

Forward(3,5)- It implies 5 year forward rate three years from now.

Q2) C) Contraction risk

Explanation - Due to decrease in interest rate people will prepay their loan to re finance it at a lower rate. This leads to increase in the contractual risk because the prepayment can only be reinvested at a lower rate.

Q3) C) 114%

Explanation -

Realised return = final value - initial value / initial value.  

= 210.59 - 98.42 / 98.42

= 112.17 / 98.42

= 1.139 ×100

= 113.9% or 114%